a ship leaves port at 1 pm traveling north at the speed of 30 miles per hour. At 3 pm the ship adjust its course 20 degrees eastward, how far was the ship from the port at 4pm. What law we will use to answer the problem


1
Expert's answer
2021-07-13T11:21:01-0400
"c=(3-1)\cdot30=60(mi)"

"a=(4-3)\cdot30=30(mi)"

"\beta=20\degree"

The Law of Cosines


"b^2=a^2+c^2-2ac\cos(180\degree-\beta)"

"b=\sqrt{30^2+60^2+2(30)(60)\cos20\degree}\approx88.786 (mi)"